OpenStudy (anonymous):

Find the radius of a cone given the surface area and slant height. Round to the nearest tenth. Surface Area = 359 ft^2 Slant height = 15 ft

OpenStudy (anonymous):

\[Surface \space Area = (\pi)(r)(l) \] l-->slant height r-->radius Now try :D

OpenStudy (kropot72):

Surface area of a cone is given by the following formula:\[\pi rl+\pi r ^{2}\]

OpenStudy (anonymous):

@kropot72 Total surface area of cone is given by \[\pi rl+\pi r ^{2}\] \[Surface \space Area = (\pi)(r)(l)\]

OpenStudy (kropot72):

Point taken. The question does not mention the area of the base. So by implication the base area can be considered to be excluded.

Directrix (directrix):

Surface area includes the area of the base. Otherwise, the question would have specified lateral area.

OpenStudy (kropot72):

Thank you for clarification Directrix.

OpenStudy (anonymous):

Thanks Directrix for clarifying